# When Gaussian Meets Surfel: Ultra-fast High-fidelity Radiance Field Rendering

### v1

@@ 4.2. Gaussian-Surfel Joint Optimization
The Gaussians are initialized with the positions of the surfels with disgarded at the 10K-th iteration during surfel optimization. The other Gaussian properties are initialized the same way as in (Kerbl et al., 2023). We periodically densify and prune 3D Gaussians during joint optimization. Every 1000 iterations, we compute the squared error map between the ground truth and the rendered image across all training views. The error in each pixel is normalized in the error map, and we treat the normalized error as the sampling probability for the pixel. We then sample a fixed number of pixels based on the probabilities for each image. Using the corresponding surfel depth map and camera parameters, we obtain a point cloud, which serves as the initial positions for the newly added 3D Gaussians. The other Gaussian properties are also initialized the same way as in (Kerbl et al., 2023). For Gaussian pruning, similar to RadSplat (Niemeyer et al., 2024), we compute the contribution score for each 3D Gaussian in the -th training view using
@@ 6.1. Mip-GES
To reduce aliasing artifacts, we utilize MSAA in surfel rendering, and employ a screen-space fixed-size EWA filter when rasterizing 3D Gaussians. However, the fixed-size EWA filter shows dilation or high-frequency artifacts when rendering at varying scales. Therefore, we apply the world space filter and the approximated screen space box filter proposed by MipSplat (2024b) to our 3D Gaussians, which can significantly reduce aliasing artifacts in renderred images.
@@ 6.2. Speedy-GES
Our surfel pruning strategy effectively removes more than 80% of invisible or low-coverage surfels, providing reliable and compact geometry representation. Although our Gaussian pruning strategy also removes low-contribution Gaussians, a considerable amount of redundant Gaussians can be further pruned at the cost of a slight quality loss. In Speedy-GES, we replace our contribution score used in Gaussian pruning by the Hessian pruning score proposed in SpeedySplat (2024), resulting in acceleration in rendering.
@@ 6.3. Compact-GES
We use the storage compression method in C3DGS (Lee et al., 2024) to further compact the basic GES. Specifically, we replace the SH coefficients of both surfels and Gaussians by querying colors from the hash grid and use residual vector quantization to quantize the scaling and rotation parameters of both surfels and Gaussians, achieving over 20 times storage compression with only a little quality loss.
@@
where and are the depth and normal of the Gaussian covering the corresponding pixel . We follow previous works (Guédon and Lepetit, 2023; Ye et al., 2024a) to use the central depth and minimal axis direction of a 3D Gaussian as and , respectively, which can effectively improve the geometry quality. However, as mentioned in 2DGS (Huang et al., 2024), simply using the central depth can not ensure the geometry consistency in different views. Therefore, we further propose 2D-GES, which replaces 3D Gaussians with 2D Gaussians, to further improve the geometry reconstruction. The and are replaced by the planar depth and normal of the 2D Gaussian in 2D-GES. We also design a specific optimization algorithm for 2D-GES with geometry regularization terms. For further details, please refer to the supplementary material.
@@ 7. Results and Evaluation
Our evaluation uses the same datasets as in (Kerbl et al., 2023), which includes eight synthetic scenes from the NeRF Synthetic Dataset (Mildenhall et al., 2020), nine real scenes from the Mip-NeRF360 Dataset (Barron et al., 2022b), two real scenes from the DeepBlending Dataset (Hedman et al., 2018b) and two real scenes from the Tanks & Temples Dataset (Knapitsch et al., 2017). These scenes include synthetic objects with complex geometries as well as various texture-rich indoor and outdoor environments, suitable for a comprehensive evaluation of different methods. Following 3DGS, we use the pre-downscaled images for the training and evaluation on real scenes. We also use the DTU Dataset (Jensen et al., 2014), which comprises 15 real scenes with corresponding masks and ground truth point clouds, to evaluate the geometry reconstruction quality of 2D-GES.
@@
We compare our basic GES representation (named 3D-GES) and its extended versions (Mip-GES, Speedy-GES, Compact-GES, 2D-GES) against the following state-of-the-art baselines: 3DGS (Kerbl et al., 2023), the vanilla 3D Gaussian Splatting; 2DGS (Huang et al., 2024), a method using 2D Gaussians as primitives to improve geometry reconstruction; SortFreeGS (Hou et al., 2024b), a method replacing the alpha blending of Gaussians with the weighted sum of Gaussians; AdrGS (Wang et al., 2024), a method using more precise bounding boxes for splatted Gaussians to accelerate rendering; SpeedySplat (Hanson et al., 2024), a method applying a new pruning score to prune about 90% of Gaussians to accelerate rendering; MipSplat (Yu et al., 2024a), a method using a 3D world space filter and a 2D screen space filter to achieve alias-free rendering; AbsGS (Ye et al., 2024b), a method using homodirectional view-space positional gradient as the criterion for densification; C3DGS (Lee et al., 2024), a method focusing on reducing the storage overhead; StopThePop (Radl et al., 2024), a method using a hierarchical sorting strategy to alleviate popping artifacts; as well as three NeRF-related methods: MipNeRF (Barron et al., 2022b), INGP (Müller et al., 2022) and Plenoxel (Fridovich-Keil et al., 2022). These baseline methods are the state-of-the-arts in the directions of rendering quality, rendering speed, storage overhead, geometry reconstruction, and view consistency, allowing for a comprehensive evaluation of our approach.
We use the standard PSNR, LPIPS, and SSIM metrics to evaluate the rendering quality. We use FPS, computed by the reciprocal of the average rendering time, to evaluate the rendering speed. We use the chamfer distance (CD) to evaluate the geometry reconstruction quality. We also follow StopThePop (Radl et al., 2024) to use and to evaluate the short-term and long-term popping artifacts on the camera paths we generated.
@@ 7.1. Comparisons With Baselines
As shown in Table 1, our 3D-GES and Mip-GES achieve high rendering quality competitive with the state-of-the-art methods. As shown in Fig. 5, 3D-GES reconstructs fine appearance details comparable to or even better than 3DGS. Among the rendering acceleration methods, AdrGS uses the load balancing loss to equalize the number of Gaussians covering each pixel. However, it results in smeared or blurry rendering results, as observed in the floor in Counter and the grass in Bicycle. SpeedySplat achieves significant rendering acceleration by discarding a large number of Gaussians, but this comes at the cost of missing appearance details, leading to poor LPIPS scores. Our Speedy-GES, by avoiding the sorting time, is able to retain more Gaussians to preserve high quality rendering and meanwhile possess ultra-fast rendering speed comparable to SpeedySplat. As shown in Fig. 6, in order to achieve a rendering quality comparable to our Speedy-GES, SpeedySplat needs to prune much fewer Gaussians and its frame rates drops to less than half of ours. In Fig. 7, we compare our method with AbsGS (Ye et al., 2024b). The densification strategy in AbsGS reduces image blurs, but still misses some details, and the aggressive densification also tends to generate floaters. Compared to the sorting-free method, SortFreeGS, our method leverages surfels to construct the coarse geometry, avoiding color leakage of occluded objects. As shown in Fig. 8, SortFreeGS exhibits noticeable color bleeding when the distance between the camera and objects is different from the training set. It is because SortFreeGS directly uses the Gaussian depths to compute the rendering weights of Gaussians, which cannot guarantee the occlusion correctness from arbitrary views. By applying the world space filter and screen space filter to 3D Gaussians, our Mip-GES achieves alias-free rendering, alleviating the dilation artifacts in distant background and high-frequency artifacts in thin structures, as shown in Fig. 9.
@@
Our GESs are sorting-free, fundamentally avoiding the popping artifacts caused by the sorting approximation in 3DGS. To evaluate view consistency, we follow (Hou et al., 2024b) to slightly rotate the camera and show two adjacent frames in Fig. 12. Please also refer to our supplementary video for a more pronounced comparison. We can see that Both 3D-GES and 2D-GES have no popping artifacts, while 3DGS, StopThePop and SpeedySplat all show popping artifacts, especially for SpeedySplat, because it removes a large number of Gaussians, resulting in a higher overall opacity of the remaining Gaussians, which exacerbates popping artifacts. StopThePop cannot guarantee to eliminate the popping artifacts, especially when the view is far from the training views. Table 2 shows the short-term and long-term metrics, which measure the consistency between rendered frames and warped frames with optical flow (Radl et al., 2024). As shown, our sort-free rendering effectively eliminates the popping artifacts and achieves the best results. Please note that as mentioned in 2DGS (Huang et al., 2024), 3DGS utilizes view-dependent intersection planes of 3D Gaussians for splatting, which also results in inconsistency. For the same reason, our 3D-GES with 3D Gaussians cannot fully guarantee the view consistency, while our 2D-GES, using 2D surfels and 2D Gaussians as primitives, achieves complete view consistency. Nevertheless, in our view consistency evaluation, we do not observe obvious view inconsistency artifacts in 3D-GES rendering.
@@ 7.2. Ablation Study
In our 2D-GES implementation, 2D Gaussians not only enrich appearance details but also play a role in smoothing the geometry. As shown in Fig. 10, the normal map rendered from surfels exhibits discontinuities, whereas 2D Gaussians effectively smooth the discontinuous areas across surfels. As shown in Table 4, 2D-GES achieves comparable geometry reconstruction quality as 2DGS on the DTU dataset. Moreover, the opaque surfels is geometrically consistent across different views, making 2D-GES capable of well reconstructing glossy surfaces, as shown in Fig. 11, while 2DGS has hole artifacts on glossy areas. It is worth noting that if we follow previous methods (Guédon and Lepetit, 2023; Ye et al., 2024a) and use the central depth and the shortest axis direction of the 3D Gaussian as the depth and normal for each point on the Gaussian, the geometry can also be optimized using the same optimization process as 2D-GES. This approach achieves better geometry compared to 3DGS, as shown in Table 4.
@@
To validate the necessity of our two-stage optimization, we experimented with a single-stage optimization scheme. Specifically, we use translucent surfels to render both depth map and color map as in 2DGS (Huang et al., 2024). The Gaussians then accumulate colors and weights according to the depth map, and the final color is computed using Eq. (5). We gradually increase to make all surfels fully opaque, identical to Sec. 4.1. As demonstrated in the original 3DGS paper, the 3D Gaussian primitives are highly flexible and can fit input images well alone. The single-stage optimization tends to favor 3D Gaussians for image fitting rather than surfels. As shown in Fig. 13, without sufficient surfels covering the scene surface, color leakage often happens in novel views, because our Gaussian primitives are blended order-independently. Quantitatively, as shown in Table 5, the averaged PSNR of the single-stage optimization on Mip-NeRF360 dataset is 27.04, while the two-stage optimization achieves 27.38, which also shows the necessity of our two-stage optimization design.
@@ 8. Discussion and Conclusion
The training time for 3D-GES is approximately 1.3 to 1.6 longer than that of 3DGS, mainly due to the time required for surfel optimization. Additionally, as shown in Fig. 17, for specular surfaces, 3DGS can rely on a large number of low-opacity Gaussians to mimic the reflection effect, but our opaque surfels lack this capability, leading to a decrease in rendering quality. Using ray tracing (Moenne-Loccoz et al., 2024b) or environment maps (Ye et al., 2024a) to capture reflections might help address this issue. Moreover, our methods and 3DGS share the same problem of initialization sensitivity. The reconstruction quality degrades when using randomly initialized points instead of SfM points in real scenes. Incorporating the deterministic state transition of MCMC samples proposed by 3DGS-MCMC (Kheradmand et al., 2024) may improve reconstruction quality when using randomly initialized points.
@@
Acknowledgements
This work is partially supported by NSF China (No. 62421003, 62227806 & 62322209) and the XPLORER PRIZE.
Appendix
@@ Appendix A Implementation details of 2D-GES
We use the same normal consistency loss as in (Huang et al., 2024) in the Gaussian-Surfel joint optimization stage, which aligns with the surface normals derived from the gradient of . To ensure that the 2D Gaussians are distributed around surfels, we introduce
@@
In the surfel optimization stage, we also apply a similar loss to align the surfel rendered normal map with the normals from surfel rendered depth map. Note only performs well during the early iterations on translucent surfels. In later iterations, has no effect on the opaque regions of the surfels with large , because the depth gradients and normals in opaque regions are naturally consistent. To obtain the accurate surfel depth map and normal map as a good basic geometry, we use the alpha-blended normal map rendered at the midpoint of optimization to supervise the alpha-blended depth and normal maps rendered in later iterations. Specifically, we render the alpha-blended depth and normal maps using the equation following 2DGS (Huang et al., 2024)
@@
We also implemented an OpenGL version of 2D-GES, which differs from 3D-GES only in that the second pass generates 2D Gaussians from points. It is worth noting that since the 2D Gaussian depth uses the planar depth rather than the center depth and the perspective projection is accurate (Huang et al., 2024), it is not suitable to use existing screen space filters for anti-aliasing in per-splat rendering. Therefore, we back-project the screen space EWA filter into the object space (Ren et al., 2002b), converting the convolution into an adjustment of the 2D Gaussian scaling for anti-aliasing, which makes our shader implementation concise and efficient. Specifically, we denote a 2D Gaussian function as , where the subscript is its covariance. The filtered 2D Gaussian function in screen space is
where is the affine approximation of the object space to screen space transformation matrix, is the EWA filter size, and is the pixel position and the rasterized position of Gaussian center. Following (Ren et al., 2002b), we back-project the screen space filter into the local Gaussian space (object space), yielding the filtered 2D Gaussian function in object space

### v2

@@ 4.2. Gaussian-Surfel Joint Optimization
The Gaussians are initialized with the positions of the surfels with disgarded at the 10K-th iteration during surfel optimization. The other Gaussian properties are initialized the same way as in (3DGS). We periodically densify and prune 3D Gaussians during joint optimization. Every 1000 iterations, we compute the squared error map between the ground truth and the rendered image across all training views. The error in each pixel is normalized in the error map, and we treat the normalized error as the sampling probability for the pixel. We then sample a fixed number of pixels based on the probabilities for each image. Using the corresponding surfel depth map and camera parameters, we obtain a point cloud, which serves as the initial positions for the newly added 3D Gaussians. The other Gaussian properties are also initialized the same way as in (3DGS). For Gaussian pruning, similar to RadSplat (niemeyer2024radsplat), we compute the contribution score for each 3D Gaussian in the -th training view using
@@ 6.1. Mip-GES
To reduce aliasing artifacts, we utilize MSAA in surfel rendering, and employ a screen-space fixed-size EWA filter when rasterizing 3D Gaussians. However, the fixed-size EWA filter shows dilation or high-frequency artifacts when rendering at varying scales. Therefore, we apply the world space filter and the approximated screen space box filter proposed by MipSplat (yu2024mip) to our 3D Gaussians, which can significantly reduce aliasing artifacts in renderred images.
@@ 6.2. Speedy-GES
Our surfel pruning strategy effectively removes more than 80% of invisible or low-coverage surfels, providing reliable and compact geometry representation. Although our Gaussian pruning strategy also removes low-contribution Gaussians, a considerable amount of redundant Gaussians can be further pruned at the cost of a slight quality loss. In Speedy-GES, we replace our contribution score used in Gaussian pruning by the Hessian pruning score proposed in SpeedySplat (hanson2024speedy), resulting in acceleration in rendering.
@@ 6.3. Compact-GES
We use the storage compression method in C3DGS (lee2024compact) to further compact the basic GES. Specifically, we replace the SH coefficients of both surfels and Gaussians by querying colors from the hash grid and use residual vector quantization to quantize the scaling and rotation parameters of both surfels and Gaussians, achieving over 20 times storage compression with only a little quality loss.
@@
where and are the depth and normal of the Gaussian covering the corresponding pixel . We follow previous works (guedon2023sugar; deferred3dgs) to use the central depth and minimal axis direction of a 3D Gaussian as and , respectively, which can effectively improve the geometry quality. However, as mentioned in 2DGS (2dgs2024), simply using the central depth can not ensure the geometry consistency in different views. Therefore, we further propose 2D-GES, which replaces 3D Gaussians with 2D Gaussians, to further improve the geometry reconstruction. The and are replaced by the planar depth and normal of the 2D Gaussian in 2D-GES. We also design a specific optimization algorithm for 2D-GES with geometry regularization terms. For further details, please refer to the supplementary material.
@@ 7. Results and Evaluation
Our evaluation uses the same datasets as in (3DGS), which includes eight synthetic scenes from the NeRF Synthetic Dataset (nerf), nine real scenes from the Mip-NeRF360 Dataset (mip_nerf_360), two real scenes from the DeepBlending Dataset (hedman2018deep) and two real scenes from the Tanks & Temples Dataset (knapitsch2017tanks). These scenes include synthetic objects with complex geometries as well as various texture-rich indoor and outdoor environments, suitable for a comprehensive evaluation of different methods. Following 3DGS, we use the pre-downscaled images for the training and evaluation on real scenes. We also use the DTU Dataset (jensen2014large), which comprises 15 real scenes with corresponding masks and ground truth point clouds, to evaluate the geometry reconstruction quality of 2D-GES.
@@
We compare our basic GES representation (named 3D-GES) and its extended versions (Mip-GES, Speedy-GES, Compact-GES, 2D-GES) against the following state-of-the-art baselines: 3DGS (3DGS), the vanilla 3D Gaussian Splatting; 2DGS (2dgs2024), a method using 2D Gaussians as primitives to improve geometry reconstruction; SortFreeGS (hou2024sortfree), a method replacing the alpha blending of Gaussians with the weighted sum of Gaussians; AdrGS (adr2024), a method using more precise bounding boxes for splatted Gaussians to accelerate rendering; SpeedySplat (hanson2024speedy), a method applying a new pruning score to prune about 90% of Gaussians to accelerate rendering; MipSplat (mipgs2024), a method using a 3D world space filter and a 2D screen space filter to achieve alias-free rendering; AbsGS (ye2024absgs), a method using homodirectional view-space positional gradient as the criterion for densification; C3DGS (lee2024compact), a method focusing on reducing the storage overhead; StopThePop (stopthepop2024), a method using a hierarchical sorting strategy to alleviate popping artifacts; as well as three NeRF-related methods: MipNeRF (mip_nerf_360), INGP (instant_ngp) and Plenoxel (Plenoxels2022). These baseline methods are the state-of-the-arts in the directions of rendering quality, rendering speed, storage overhead, geometry reconstruction, and view consistency, allowing for a comprehensive evaluation of our approach.
We use the standard PSNR, LPIPS, and SSIM metrics to evaluate the rendering quality. We use FPS, computed by the reciprocal of the average rendering time, to evaluate the rendering speed. We use the chamfer distance (CD) to evaluate the geometry reconstruction quality. We also follow StopThePop (stopthepop2024) to use and to evaluate the short-term and long-term popping artifacts on the camera paths we generated.
@@ 7.1. Comparisons With Baselines
As shown in Table 1, our 3D-GES and Mip-GES achieve high rendering quality competitive with the state-of-the-art methods. As shown in Fig. 5, 3D-GES reconstructs fine appearance details comparable to or even better than 3DGS. Among the rendering acceleration methods, AdrGS uses the load balancing loss to equalize the number of Gaussians covering each pixel. However, it results in smeared or blurry rendering results, as observed in the floor in Counter and the grass in Bicycle. SpeedySplat achieves significant rendering acceleration by discarding a large number of Gaussians, but this comes at the cost of missing appearance details, leading to poor LPIPS scores. Our Speedy-GES, by avoiding the sorting time, is able to retain more Gaussians to preserve high quality rendering and meanwhile possess ultra-fast rendering speed comparable to SpeedySplat. As shown in Fig. 6, in order to achieve a rendering quality comparable to our Speedy-GES, SpeedySplat needs to prune much fewer Gaussians and its frame rates drops to less than half of ours. In Fig. 7, we compare our method with AbsGS (ye2024absgs). The densification strategy in AbsGS reduces image blurs, but still misses some details, and the aggressive densification also tends to generate floaters. Compared to the sorting-free method, SortFreeGS, our method leverages surfels to construct the coarse geometry, avoiding color leakage of occluded objects. As shown in Fig. 8, SortFreeGS exhibits noticeable color bleeding when the distance between the camera and objects is different from the training set. It is because SortFreeGS directly uses the Gaussian depths to compute the rendering weights of Gaussians, which cannot guarantee the occlusion correctness from arbitrary views. By applying the world space filter and screen space filter to 3D Gaussians, our Mip-GES achieves alias-free rendering, alleviating the dilation artifacts in distant background and high-frequency artifacts in thin structures, as shown in Fig. 9.
@@
Our GESs are sorting-free, fundamentally avoiding the popping artifacts caused by the sorting approximation in 3DGS. To evaluate view consistency, we follow (hou2024sortfree) to slightly rotate the camera and show two adjacent frames in Fig. 12. Please also refer to our supplementary video for a more pronounced comparison. We can see that Both 3D-GES and 2D-GES have no popping artifacts, while 3DGS, StopThePop and SpeedySplat all show popping artifacts, especially for SpeedySplat, because it removes a large number of Gaussians, resulting in a higher overall opacity of the remaining Gaussians, which exacerbates popping artifacts. StopThePop cannot guarantee to eliminate the popping artifacts, especially when the view is far from the training views. Table 2 shows the short-term and long-term metrics, which measure the consistency between rendered frames and warped frames with optical flow (stopthepop2024). As shown, our sort-free rendering effectively eliminates the popping artifacts and achieves the best results. Please note that as mentioned in 2DGS (2dgs2024), 3DGS utilizes view-dependent intersection planes of 3D Gaussians for splatting, which also results in inconsistency. For the same reason, our 3D-GES with 3D Gaussians cannot fully guarantee the view consistency, while our 2D-GES, using 2D surfels and 2D Gaussians as primitives, achieves complete view consistency. Nevertheless, in our view consistency evaluation, we do not observe obvious view inconsistency artifacts in 3D-GES rendering.
@@ 7.2. Ablation Study
In our 2D-GES implementation, 2D Gaussians not only enrich appearance details but also play a role in smoothing the geometry. As shown in Fig. 10, the normal map rendered from surfels exhibits discontinuities, whereas 2D Gaussians effectively smooth the discontinuous areas across surfels. As shown in Table 4, 2D-GES achieves comparable geometry reconstruction quality as 2DGS on the DTU dataset. Moreover, the opaque surfels is geometrically consistent across different views, making 2D-GES capable of well reconstructing glossy surfaces, as shown in Fig. 11, while 2DGS has hole artifacts on glossy areas. It is worth noting that if we follow previous methods (guedon2023sugar; deferred3dgs) and use the central depth and the shortest axis direction of the 3D Gaussian as the depth and normal for each point on the Gaussian, the geometry can also be optimized using the same optimization process as 2D-GES. This approach achieves better geometry compared to 3DGS, as shown in Table 4.
@@
To validate the necessity of our two-stage optimization, we experimented with a single-stage optimization scheme. Specifically, we use translucent surfels to render both depth map and color map as in 2DGS (2dgs2024). The Gaussians then accumulate colors and weights according to the depth map, and the final color is computed using Eq. (5). We gradually increase to make all surfels fully opaque, identical to Sec. 4.1. As demonstrated in the original 3DGS paper, the 3D Gaussian primitives are highly flexible and can fit input images well alone. The single-stage optimization tends to favor 3D Gaussians for image fitting rather than surfels. As shown in Fig. 13, without sufficient surfels covering the scene surface, color leakage often happens in novel views, because our Gaussian primitives are blended order-independently. Quantitatively, as shown in Table 5, the averaged PSNR of the single-stage optimization on Mip-NeRF360 dataset is 27.04, while the two-stage optimization achieves 27.38, which also shows the necessity of our two-stage optimization design.
@@ 8. Discussion and Conclusion
The training time for 3D-GES is approximately 1.3 to 1.6 longer than that of 3DGS, mainly due to the time required for surfel optimization. Additionally, as shown in Fig. 17, for specular surfaces, 3DGS can rely on a large number of low-opacity Gaussians to mimic the reflection effect, but our opaque surfels lack this capability, leading to a decrease in rendering quality. Using ray tracing (moenne20243d) or environment maps (deferred3dgs) to capture reflections might help address this issue. Moreover, our methods and 3DGS share the same problem of initialization sensitivity. The reconstruction quality degrades when using randomly initialized points instead of SfM points in real scenes. Incorporating the deterministic state transition of MCMC samples proposed by 3DGS-MCMC (kheradmand20243d) may improve reconstruction quality when using randomly initialized points.
@@
Acknowledgements.
This work is partially supported by NSF China (No. 62421003, 62227806 & U23A20311) and the XPLORER PRIZE.Appendix
@@ Appendix A Implementation details of 2D-GES
We use the same normal consistency loss as in (2dgs2024) in the Gaussian-Surfel joint optimization stage, which aligns with the surface normals derived from the gradient of . To ensure that the 2D Gaussians are distributed around surfels, we introduce
@@
In the surfel optimization stage, we also apply a similar loss to align the surfel rendered normal map with the normals from surfel rendered depth map. Note only performs well during the early iterations on translucent surfels. In later iterations, has no effect on the opaque regions of the surfels with large , because the depth gradients and normals in opaque regions are naturally consistent. To obtain the accurate surfel depth map and normal map as a good basic geometry, we use the alpha-blended normal map rendered at the midpoint of optimization to supervise the alpha-blended depth and normal maps rendered in later iterations. Specifically, we render the alpha-blended depth and normal maps using the equation following 2DGS (2dgs2024)
@@
We also implemented an OpenGL version of 2D-GES, which differs from 3D-GES only in that the second pass generates 2D Gaussians from points. It is worth noting that since the 2D Gaussian depth uses the planar depth rather than the center depth and the perspective projection is accurate (2dgs2024), it is not suitable to use existing screen space filters for anti-aliasing in per-splat rendering. Therefore, we back-project the screen space EWA filter into the object space (ren2002object), converting the convolution into an adjustment of the 2D Gaussian scaling for anti-aliasing, which makes our shader implementation concise and efficient. Specifically, we denote a 2D Gaussian function as , where the subscript is its covariance. The filtered 2D Gaussian function in screen space is
where is the affine approximation of the object space to screen space transformation matrix, is the EWA filter size, and is the pixel position and the rasterized position of Gaussian center. Following (ren2002object), we back-project the screen space filter into the local Gaussian space (object space), yielding the filtered 2D Gaussian function in object space